# Bound States in Lee’s Complex Ghost Model

### v1

@@ Abstract
Quantum field theories (QFTs) including fourth-derivative terms such as the Lee-Wick finite QED and quadratic gravity have a better ultra-violet behavior compared to standard theories with second-derivative ones, but the existence of ghost with negative norm endangers unitarity. Such a ghost in general acquires a pair of complex conjugate masses from radiative corrections whose features are concisely described by the so-called Lee model. Working with the canonical operator formalism of QFTs, we investigate the issue of bound states in the Lee model. We find that the bound states cannot be created from ghosts by contributions of a complex delta function, which is a complex generalization of the well-known Dirac delta function. Since the cause of unitarity violation in the Lee-Wick model is the existence of the complex delta function instead of the Dirac delta function, it is of interest to notice that the violation of the unitarity is also connected to the non-existence of bound states. Finally, the problem of amelioration of the unitarity in quadratic gravity is briefly discussed.
@@ 3 Correlation function of composite operator
where we have used Eq. (3.5) in the last line. This equation implies that if there were a pole at some on-shell value in Eq. (LABEL:C(p)), $p^{2}=-m^{2}$, we would have a bound state with the mass $m$. Thus, by solving this equation in the next section, we will investigate the problem of whether there is such a bound state or not.
@@ 4 Bound states
Now we are ready to calculate the first term on the RHS of Eq. (4.12). This integral is a usual one-loop integral except for complex masses. First, we would like to make the Wick rotation, by which the $k^{0}$ integration contour along the real axis is rotated to the imaginary axis. In contrast to the conventional situation, the contour now passes over two poles at $k^{0}=\pm\omega_{k}$. [^6] In addition to the two poles, the contour might pass over a pole at $k^{0}=\sqrt{\vec{p}\,^{2}+m^{2}}-\omega_{p-k}^{*}$. Here we assume ${\rm{Re}}(\sqrt{\vec{p}\,^{2}+m^{2}}-\omega_{p-k}^{*})<0$, which depends on the magnitudes of $m,{\rm{Re}}\,M^{2}$ and $|\vec{k}|$. If not, we would get an additional residue at this pole, but our final conclusion remains unchanged. The evaluation of contributions from the poles leads to
@@
$=$ $\int_{I}d^{4}k\,\frac{1}{k^{2}+M^{2}}\frac{1}{(p-k)^{2}+M^{*2}}-\int d^{3}k\,\frac{\pi i}{\omega_{k}}\sum_{\pm}\frac{1}{-(p^{0}\mp\omega_{k})^{2}+(\omega_{p-k}^{*})^{2}},$
where $I$ denotes the imaginary axis running from $(0,-i\infty)$ to $(0,i\infty)$ in the complex $k^{0}$-plane. Note that the second term just cancels the second term of Eq. (4.12). To put it differently, we can rewrite the first and second terms of Eq. (4.12) into the integral over the imaginary axis [^7] In Ref. [5], the authors start with the Euclidean Lagrangian density and calculate only the same quantity as $K(p)$ on the basis of path integral approach.
$K(p)$ $\equiv$ $\int_{R}d^{4}k\,\frac{1}{k^{2}+M^{2}}\frac{1}{(p-k)^{2}+M^{*2}}+\int d^{3}k\,\frac{\pi i}{\omega_{k}}\sum_{\pm}\frac{1}{-(p^{0}\mp\omega_{k})^{2}+(\omega_{p-k}^{*})^{2}}$ (4.15)
$=$ $\int_{I}d^{4}k\,\frac{1}{k^{2}+M^{2}}\frac{1}{(p-k)^{2}+M^{*2}}.$
Next it is straightforward to calculate $K$. We introduce $k^{4}$ defined by $k^{0}=ik^{4}$. Then we have $k^{2}=-(k^{0})^{2}+\vec{k}^{2}=(k^{4})^{2}+\vec{k}^{2}\equiv k_{E}^{2}$. Using these Euclidean momenta, $K(p)$ can be recast to the form
$K(p)=i\int d^{4}k_{E}\,\frac{1}{k_{E}^{2}+M^{2}}\frac{1}{(p_{E}-k_{E})^{2}+M^{*2}},$ (4.16)
where we have introduced the Euclidean momenta $p_{E}$ as well. Below we will write $k,p$ instead of $k_{E},p_{E}$ for convenience. By means of the Feyman’s parameter formula, $K(p)$ can be rewritten as follows:
$K(p)$ $=$ $i\int d^{4}k\int_{0}^{1}dx\,\frac{1}{\{[k-p(1-x)]^{2}+p^{2}x(1-x)+M^{*2}+(M^{2}-M^{*2})x\}^{2}}$ (4.17)
$=$ $i\int d^{4}k\int_{0}^{1}dx\,\frac{1}{[k^{2}+p^{2}x(1-x)+M^{*2}+(M^{2}-M^{*2})x]^{2}}$
$=$ $i\int_{0}^{1}dx\int d^{4}k\,\frac{1}{(k^{2}+\Delta)^{2}},$
where in the second line we have shifted the integration variable [^8] Precisely speaking, this procedure is legitimate only after regularizing the integral in a proper manner. as $k^{\mu}\rightarrow k^{\mu}+p^{\mu}(1-x)$ and in the last line we have defined
$\Delta\equiv p^{2}x(1-x)+M^{*2}+(M^{2}-M^{*2})x.$ (4.18)
It is obvious that $K(p)$ is logarithmically divergent so we adopt the Pauli-Villars regularization. To do that, let us consider the following quantity:
@@
where we have used the spherical coordinates in the second line. Thus, from Eqs. (4.17)-(4.20), $K(p)$ takes the form:
$K(p)$ $=$ $-i\pi^{2}\int_{0}^{1}dx\,\log\frac{p^{2}x(1-x)+M^{*2}+(M^{2}-M^{*2})x}{\Lambda^{2}}$ (4.21)
@@
- $\int d^{3}k\sum_{\pm}\frac{\pi^{2}}{\omega_{k}\omega_{p-k}^{*}}\delta_{c}[\pm(\omega_{p-k}^{*}+\omega_{k})-p^{0}]\Bigg)=0.$
Here let us suppose that there is no the last term on the LHS, which involves the complex delta function. Then, since the second term on the LHS including the logarithmic function, i.e., the function $K(p)$, has a rather smooth expression with respect to $p^{2}(=-m^{2})$, there could be a possibility such that we have a nontrivial solution to Eq. (4.23) as suggested in Ref. [5, 6]. However, the presence of the last term forbits such a solution. In other words, there is no non-trivial solution to Eq. (4.23) owing to the existence of the last term involving the complex delta function. It is worthwhile to stress that it is the term that breaks the unitarity of the Lee model [3, 4]. To put it differently, the term with the complex delta function not only breaks the unitarity but also prohibits the existence of the bound state made out of complex ghost fields.
Finally, let us comment on the other choices of composite operators constructed from ghost fields such as ${\cal{O}}(x)=\varphi^{\dagger 2}(x)+\varphi^{2}(x)$ etc. We have investigated several candidates of composite operators and found that the complex delta functions always appear and prohibit the existence of nontrivial solutions to the pole equation. It is therefore concluded that there is no bound state constructed from ghost fields in the Lee model.
@@ 5 Conclusions
In this paper, based on the canonical operator formalism we have demonstrated that there is no bound state in the Lee model which nicely describes some features of QFTs with fourth-derivative terms. We should stress that we have only made use of the operator formalism, which is the basic tool of QFTs.
Here it is interesting to imagine whether if there were a bound state with positive norm, it could give us a resolution to the problem of massive ghost, which violates the unitarity, in the higher derivative theories [14, 15, 16, 17, 18, 19, 20]. We can show that that is not the case from two reasons. One of the two reasons is that as mentioned in the introduction, even if there would be a bound state made from ghost fields, the bound state must be not a normal bound state with positive norm but a ghost bound state with negative norm by the conservation law of the total unitarity of the S-matrix. The other reason is that the construction of the bound state does not mean the permanent confinement of the massive ghost and the bound state is dissolved into elementary ghost fields in the weak coupling regime. To solve the problem of the massive ghost in the higher derivative theories such as quadratic gravity, we need the permanent confinement such as the confinement of quarks and gluons in QCD.
Under such a situation, it is valuable to pursue a new solution to the massive ghost problem from the viewpoint of bound states within the framework of the canonical formalism. For instance, in the analysis of physical modes of quadratic gravity [21] [^9] The BRST formalism of various gravitational theories has been already constructed in the de Donder (harmonic) gauge [22, 23, 24, 25, 26, 27, 28]., we usually assume that there is no bound state. If this assumption is not imposed, we are free to take account of a bound state which might be constructed from Faddeev-Popov (FP) ghosts and the other elementary fields. Of course, as mentioned above, such a bound state carries a negative norm so it must be a ghost bound state. However, if a certain bound state is constructed from the BRST ghost and its BRST-conjugate is also formed, they together constitute a BRST doublet, thereby belonging to unphysical state and nullifying the massive ghost problem [29] since this mechanism realizes the scenario of the permanent confinement of the massive ghost. The present formalism would be useful in clarifying the (non-)existence of such the bound state. We wish to return to this problem in the near future.

### v2

@@ Abstract
Quantum field theories (QFTs) including fourth derivative terms such as the Lee-Wick finite QED and quadratic gravity have a better ultra-violet behavior compared to standard theories with second derivative ones, but the existence of ghost with negative norm endangers unitarity. Such a ghost in general acquires a pair of complex conjugate masses from radiative corrections whose features are concisely described by the so-called Lee model. Working with the canonical operator formalism of QFTs, we investigate the issue of bound states in the Lee model. We find that the bound states cannot be created from ghosts by contributions of a complex delta function, which is a complex generalization of the well-known Dirac delta function. Since the cause of unitarity violation in the Lee-Wick model is the existence of the complex delta function instead of the Dirac delta function, it is of interest to notice that the violation of the unitarity is also connected to the non-existence of bound states. Finally, the problem of amelioration of the unitarity in quadratic gravity is briefly discussed.
@@ 3 Correlation function of composite operator
where we have used Eq. (3.5) in the last line. It is remakable that Eq. (LABEL:C(p)) is an exact result in the sense that all quantum corrections are summed over. This equation implies that if there were a pole at some on-shell value in Eq. (LABEL:C(p)), $p^{2}=-m^{2}$, we would have a bound state with the mass $m$. Thus, by solving this equation in the next section, we will investigate the problem of whether there is such a bound state or not.
@@ 4 Bound states
Now we are ready to calculate the first term on the RHS of Eq. (4.12), which is denoted as $K(p)$ from now on. Note that this integral $K(p)$ is a usual one-loop integral except for complex masses, so we may attempt to make the Wick rotation, by which the $k^{0}$ integration contour along the real axis is rotated to the imaginary axis. However, in contrast to the conventional situation, there is a possibility that the contour passes over poles. To clarify this situation, we make use of the Feyman’s parameter formula and rewrite $K(p)$ as follows:
$K(p)$ $\equiv$ $\int_{R}d^{4}k\,\frac{1}{k^{2}+M^{2}}\frac{1}{(p-k)^{2}+M^{*2}}$ (4.14)
$=$ $\int_{R}d^{4}k\int_{0}^{1}dx\,\frac{1}{\{[k-p(1-x)]^{2}+p^{2}x(1-x)+M^{*2}+(M^{2}-M^{*2})x\}^{2}}$
$=$ $\int_{0}^{1}dx\int_{R}d^{4}q\,\frac{1}{(q^{2}+\Delta)^{2}},$
where in the last line we have made the change of variables, q=kp(1x)q=k-p(1-x) and defined
$\Delta\equiv p^{2}x(1-x)+M^{*2}+(M^{2}-M^{*2})x.$ (4.15)
The poles are located at $q^{0}=\pm\sqrt{\vec{q}\,^{2}+\Delta}$ in the complex $q^{0}$-plane. The imaginary part of the poles comes from the quantity $\Delta$ and is given by
${\rm{Im}}\,\Delta=(2x-1){\rm{Im}}\,M^{2},$ (4.16)
which becomes positive for $\frac{1}{2}<x<1$ due to ${\rm{Im}}\,M^{2}>0$. In this case, since the poles are located in the first and third quadrants, we have to evaluate the residues from the two poles in making the Wick rotation. [^6] For simplicity, we assume $\vec{q}\,^{2}+{\rm{Re}}\,\Delta>0$. After all, the evaluation of contributions from the poles leads to
@@
$=$ $\int_{I}d^{4}k\,\frac{1}{k^{2}+M^{2}}\frac{1}{(p-k)^{2}+M^{*2}}+\pi i\int_{\frac{1}{2}}^{1}dx\int d^{3}q\frac{1}{(\vec{q}\,^{2}+\Delta)^{\frac{3}{2}}},$
where $I$ denotes the imaginary axis running from $(0,-i\infty)$ to $(0,i\infty)$ in the complex $k^{0}$-plane.
Next it is straightforward to calculate the first term on the RHS of Eq. (LABEL:Wick), which is now denoted as $K_{I}(p)$. [^7] In Ref. [5], the authors start with the Euclidean Lagrangian density and calculate only the same quantity as $K_{I}(p)$ on the basis of path integral approach. As in Eq. (4.14), we first rewrite $K_{I}(p)$ by means of the Feynman parameter formula and then introduce the Euclidean momentum which is defined by $q^{0}=ik^{4}$ and $\vec{q}=\vec{k}$. Then we have $q^{2}=-(q^{0})^{2}+\vec{q}\,^{2}=(k^{4})^{2}+\vec{k}^{2}\equiv k^{2}$. Using these Euclidean momenta, $K_{I}(p)$ can be recast to the form
$K_{I}(p)=i\int_{0}^{1}dx\int d^{4}k\,\frac{1}{(k^{2}+\Delta)^{2}}.$ (4.18)
It is obvious that $K_{I}(p)$ is logarithmically divergent so we adopt the Pauli-Villars regularization. To do that, let us consider the following quantity:
@@
where we have used the spherical coordinates in the second line. Thus, from Eqs. (4.15) and (4.18), $K_{I}(p)$ takes the form:
$K_{I}(p)$ $=$ $-i\pi^{2}\int_{0}^{1}dx\,\log\frac{p^{2}x(1-x)+M^{*2}+(M^{2}-M^{*2})x}{\Lambda^{2}}$ (4.21)
@@
$+$ $i\pi\int_{\frac{1}{2}}^{1}dx\int d^{3}q\frac{1}{(\vec{q}\,^{2}+\Delta)^{\frac{3}{2}}}+\int d^{3}k\frac{\pi i}{\omega_{k}}\sum_{\pm}\frac{1}{-(p^{0}\mp\omega_{k})^{2}+(\omega_{p-k}^{*})^{2}}$
- $\int d^{3}k\sum_{\pm}\frac{\pi^{2}}{\omega_{k}\omega_{p-k}^{*}}\delta_{c}[\pm(\omega_{p-k}^{*}+\omega_{k})-p^{0}]\Bigg)$
$=$ $0.$
Here let us suppose that there are no last three terms on the LHS, which involve the integration. Then, since the second term on the LHS including the logarithmic function, i.e., the function $K_{I}(p)$, has a rather smooth expression with respect to $p^{2}(=-m^{2})$, there could be a possibility such that we have a nontrivial solution to Eq. (4.23) as suggested in Ref. [5, 6]. However, the presence of the last three terms forbits such a solution. In particular, there is no non-trivial solution to Eq. (4.23) owing to the existence of the last term involving the complex delta function. It is worthwhile to stress that it is the term that breaks the unitarity of the Lee model [3, 4]. To put it differently, the term with the complex delta function not only breaks the unitarity but also prohibits the existence of the bound state made out of complex ghost fields.
Finally, let us comment on the other choices of composite operators constructed from ghost fields such as ${\cal{O}}(x)=\varphi^{\dagger 2}(x)+\varphi^{2}(x)$ and ${\cal{O}}(x)=\varphi^{\dagger 2}(x)\varphi^{2}(x)$ etc. We have investigated several candidates of composite operators and found that the complex delta functions always appear and prohibit the existence of nontrivial solutions to the pole equation. It is therefore concluded that there is no bound state constructed from ghost fields in the Lee model.
@@ 5 Conclusions
In this paper, based on the canonical operator formalism we have demonstrated that there is no bound state in the Lee model which nicely describes some features of QFTs with fourth derivative terms. We should stress that we have only made use of the operator formalism, which is the basic tool of QFTs.
Here it is interesting to imagine whether if there were a bound state with positive norm, it could give us a resolution to the problem of massive ghost, which violates the unitarity, in the higher derivative theories [14, 15, 16, 17, 18, 19, 20, 21]. We can show that that is not the case from two reasons. One of the two reasons is that as mentioned in the introduction, even if there would be a bound state made from ghost fields, the bound state must be not a normal bound state with positive norm but a ghost bound state with negative norm by the conservation law of the total unitarity of the S-matrix. The other reason is that the construction of the bound state does not mean the permanent confinement of the massive ghost and the bound state is dissolved into elementary ghost fields in the weak coupling regime. To solve the problem of the massive ghost in the higher derivative theories such as quadratic gravity, we need the permanent confinement like the confinement of quarks and gluons in QCD.
Under such a situation, it is valuable to pursue a new solution to the massive ghost problem from the viewpoint of bound states within the framework of the canonical formalism. For instance, in the analysis of physical modes of quadratic gravity [22] [^8] The BRST formalism of various gravitational theories has been already constructed in the de Donder (harmonic) gauge [23, 24, 25, 26, 27, 28, 29]., we usually assume that there is no bound state. If this assumption is not imposed, we are free to take account of a bound state which might be constructed from Faddeev-Popov (FP) ghosts and the other elementary fields. Of course, as mentioned above, such a bound state carries a negative norm so it must be a ghost bound state. However, if a certain bound state is constructed from the massive ghost and BRST FP ghost, and its BRST-conjugate corrsponding to this bound state is also formed, they together constitute a BRST doublet, thereby belonging to unphysical state and nullifying the massive ghost problem [30] since this mechanism realizes the scenario of the permanent confinement of the massive ghost. Then, an important lesson obtained from the present article is that we should start with a higher derivative theory without complex masses and investigate the bound state problem involving the massive ghost. The present formalism is also useful in clarifying the (non-)existence of such the bound states. We wish to return to this problem in the near future.